OpenStudy (misssunshinexxoxo):

Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (misssunshinexxoxo):

Please show steps with how to get the answer

OpenStudy (anonymous):

\[y-y_1 = m(x-x_1)\] point-slope form \[y = mx+b ---> slope-intercept form \] \[ax+by = c ----> standard form \]

OpenStudy (anonymous):

the m is the slope, and to find the slope, you need the slope formula of \[\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

after you find the slope, the rest is plugging the points

OpenStudy (misssunshinexxoxo):

Please show me how to plug it in for this problem to understand better

OpenStudy (anonymous):

i take that back, the slope is already given, so you just need to plug it in

OpenStudy (misssunshinexxoxo):

What do the answers come out to?

OpenStudy (anonymous):

for the point slope form \[\huge y+2=2(x-3)\]

OpenStudy (anonymous):

now you have y+2 = 2(x-3) solve for y, and you would have it in slope intercept form which is y= mx+b

OpenStudy (misssunshinexxoxo):

I will write a review for your help and gave a medal . I have been having so much trouble with inequalities. I am a very great student and strive my best but I have trouble

OpenStudy (misssunshinexxoxo):

Please give how to get standard form once I plug it in?

OpenStudy (anonymous):